# Chaos Engineering: A Multi-Vocal Literature Review

### v1

@@ 5.1.4. Deployment
5.1.5. Evaluation Approaches
@@
Evaluation refers to determining the effectiveness of chaos engineering experiments by measuring their impact on system stability, the system’s ability to handle induced failures, and the growth and maturity of the organization’s chaos engineering practices. Evaluation approaches can be categorized into qualitative and quantitative methods.
@@
  • Business Impact Evaluation assesses how chaos experiments affect key business metrics, including revenue, customer churn, and operational efficiency (Inc., 2023; Jernberg et al., 2020). This involves understanding specific financial losses, customer complaints, and service-level breaches resulting from experiments(National Australia Bank, 2020; Palani and Gupta, 2023). For instance, a chaos experiment that causes downtime during a product launch could result in significant lost sales and a spike in customer service tickets (Ji et al., 2023; Starr et al., 2022). This evaluation is critical to linking technical failures with direct business consequences, enabling more targeted investment in stability measures where it matters most (Manna, 2021; Moon, 2022).
@@
5.2. Chaos Engineering Tools
We used the proposed taxonomy to evaluate 10 popular chaos engineering tools (T1-T10 in Table 10). We selected these tools by identifying the most-cited tools from our pool of reviewed papers, then ranking them using GitHub metrics, precisely the ’Most stars’ measure, to confirm their popularity and community engagement (Borges et al., 2016). Collectively, these tools had over 93,000 stars as of November 2024. These tools have been mapped according to their applicability to different execution environments, automation modes, strategies, and deployment stages. This categorization enables practitioners to make well-informed choices by aligning tool selection with specific infrastructure needs, operational requirements, automation capabilities, and system risk tolerance. Consequently, selecting the appropriate tool in a software development project becomes a crucial decision (Kavaler et al., 2019). For example, T1 and T2 provide manual and semi-automated chaos testing capabilities, supporting Kubernetes, Docker, and virtual machine setups. While T1 supports performing chaos experiments in live production environments, T2 focuses on testing environments, particularly pre-production setups. T3 is a retired tool designed initially for AWS but lacks support for Kubernetes. Its functions have been migrated to newer Netflix tools such as Chaos Monkey, Swabbie, and Spinnaker, making it better suited to traditional infrastructures. T4 and T5 both support Kubernetes and virtual machine environments in hybrid clouds. These tools offer full automation options, with features like workflow automation and custom integrations, making them ideal for chaos testing in pre-production and production environments. T6 and T7 focus on Kubernetes and serverless functions, offering manual, semi-automated, and fully automated modes. While T6 is optimized for hybrid cloud setups with advanced workflow customization, T7 is better suited for production resilience testing in Kubernetes environments. T8 and T9 emphasize adaptability across diverse infrastructures. T8 supports Kubernetes, serverless functions, and on-premise setups, focusing on manual testing modes with customizable workflows. T9, on the other hand, operates across virtual machines and Kubernetes environments, providing semi-automated and fully automated capabilities with workflow automation and custom integration options, making it suitable for public cloud systems. Finally, T10 excels in hybrid cloud environments, supporting Kubernetes and virtual machines with advanced automation and integration capabilities for pre-production and production stages. This taxonomy ensures that chaos engineering tools are mapped to their ideal use cases, helping practitioners choose the best tools for their systems. Our online appendix provides the complete list of tools considered.
Figure 9. A Taxonomy of Chaos Engineering Platforms.
[…]
A taxonomy of chaos engineering platforms.
Table 10. Top-10 Most Used Chaos Engineering Platforms.
Docker Kubernetes
Virtual
Machine
Serverless
Function
Private Public Hybrid
T1 Chaos Monkey
T2 Toxiproxy
T3 Simian Army
T4 Chaos Mesh
T5 Chaosblade
T6 Litmus
T7 Pumba
T8 Chaos Toolkit
T9 Powerful seal
T10 Chaoskube
5.3. Adoption Practices
Adopting chaos engineering requires a structured approach, balancing best practices with avoiding common mistakes. To effectively implement chaos experiments and reduce the risk of system failures or disruptions, organizations should follow well-established guidelines for planning, executing, and evaluating these experiments  (Siwach et al., 2022; Green, 2023). Table 11 summarizes essential best practices, such as starting with small, controlled experiments, setting clear objectives, automating tests, and closely monitoring system behavior. Additionally, organizations should limit the scope of experiments, encourage cross-team collaboration, and continuously refine processes based on experiment outcomes. While following best practices is crucial, organizations must also avoid common pitfalls, such as those outlined in Table 12, which include failing to tailor experiments to specific systems, neglecting proper planning and hypothesis formation, and lacking adequate safety mechanisms. By adhering to best practices and avoiding these common pitfalls, organizations can confidently integrate chaos engineering into their systems and achieve successful outcomes (Siwach et al., 2022).
Table 11. Best Practices
No Best Practices Description Articles Total
1 Start Small and Scale Up Start with manageable tests, then expand gradually. (Fogli et al., 2023; Torkura et al., 2020; Bremmers, 2021; Patel, 2022; Gill, 2021; Chen et al., 2022; Siwach et al., 2022; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Palani and Gupta, 2023; Kadikar, 2023; Green, 2023) 12
6 Document and Improve Record experiments to continually improve the system. (Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Siwach et al., 2022; Palani and Gupta, 2023; Kadikar, 2023; Green, 2023; Kesim et al., 2020) 11
Table 12. Bad Practices
No Bad Practices Description Articles Total
1 Misaligned Chaos Chaos experiments that don’t match the system yield misleading results. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Vu et al., 2022; Nikolaidis et al., 2023; Ma et al., 2023; Sondhi et al., 2021; Jernberg et al., 2020; Konstantinou et al., 2021; Park et al., 2019; Zhang et al., 2023; Rivera et al., 2023; Patel, 2022; Le, 2022; Lardo, 2019) 16
2 Inadequate Experimentation Without planning, clear goals, and baseline tracking, assessing chaos test impacts becomes difficult. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Jakkaraju, 2020; Le, 2022; Hawkins, 2020; Anwar and Arunachalam, 2019; Dua, 2024; National Australia Bank, 2020; Mooney, 2023; Starr et al., 2022; Palani and Gupta, 2023; Gill, 2022; Fawcett, 2020; Gianchandani et al., 2022; Bairyev, 2023; Abdul, 2024; Azure Readiness, 2023; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Frank et al., 2021b; Jernberg et al., 2020; Konstantinou et al., 2021; Ji et al., 2023; Soldani and Brogi, 2021; Park et al., 2019; Wachtel, 2023) 31
3 Safety Mechanism Oversight Not testing rollback and safety features risks uncontrolled failures. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Bhaskar, 2022; Katirtzis, 2022; Dua, 2024; Starr et al., 2022; Tivey et al., 2019; Palani and Gupta, 2023; Bairyev, 2023; Sbiai et al., 2023; Kadikar, 2023; Green, 2023; Vinisky, 2024; Pierce et al., 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Soldani and Brogi, 2021; Konstantinou et al., 2021) 25
@@ 6. Current State of Chaos Engineering (RQ4)
This section first presents the classification of research in chaos engineering, then examines the number of sources published per year and venue type, and finally reviews the venue rankings and key contributors in the field.
6.1. Number of Sources by Research Type
@@
To better understand the current state of chaos engineering research, we classified the primary studies based on the research classification framework proposed by Wieringa et al. (Wieringa et al., 2006). This framework includes the following categories:
  • Evaluation research: These studies investigate problems in software engineering practice, often uncovering new knowledge about causal or logical relationships among different phenomena. This type of research is conducted through case or field studies, field experiments, and surveys.
  • Philosophical papers: These studies introduce new conceptual frameworks or ways of thinking.
  • Solution proposal papers: These studies suggest techniques and argue for their relevance without full-scale validation.
  • Validation research: These studies rigorously examine the properties of a proposed solution that has yet to be implemented in practice, employing methodologically sound approaches such as experiments and simulations.
  • Opinion papers: These studies present the author’s views on what is good or bad about a particular issue and how things should be approached.
  • Personal experience papers: These studies describe what was done based on the author’s own experience with one or more projects rather than explaining why it was done.
@@
The first author did the classification, which was then reviewed by the second author, with all discrepancies resolved through discussion. The results of this classification, shown in Figure 12, indicate that solution proposal papers were the most frequent type of research between 2019 and 2024. In the academic literature, solution proposal papers dominated, with nine papers in 2021 and 2022 and slightly fewer (eight) in 2023. The grey literature, on the other hand, showed a marked rise in experience reports, growing from six reports in 2022 to eight in 2023. This classification highlights research activities’ evolving distribution and focus, emphasizing the dominance of solution proposals in academic literature and the increasing role of practical experience reports in grey literature.
6.2. Number of Sources by Year and Venue Type
We plot the distribution of chaos engineering studies from 2019 to 2024 to examine its adoption pattern (see Figure 10). In 2019, academic publications were minimal, limited to conferences and workshops, while grey literature focused on blogs, reflecting an early exploratory phase. By 2020, academic interest grew, focusing on conferences, though grey literature centered around blogs. In 2021, academic publications peaked across journals, conferences, and workshops, marking the “early adopter” phase, while grey literature remained stable, illustrating the challenges of Moore’s Chasm (see Figure 1). In 2022, academic output slightly declined. Still, grey literature grew, diversifying with blogs and white papers, suggesting chaos engineering was beginning to bridge Moore’s Chasm and gain broader acceptance. By 2023, academic publications remained steady, though workshops declined, while grey literature expanded further, especially with white papers, signaling continued industry engagement. The slight decline observed in 2024 for both academic and grey literature suggests a possible temporary slowdown, as the year is not yet complete, and more publications may emerge (Werbińska-Wojciechowska and Winiarska, 2023).
A precise distribution emerges between academic and grey literature publications when examining venue types. Of the 44 scholarly publications, 28 (63.64%) were conference papers, 11 (25%) were journal articles, and 5 (11.36%) were workshops, reflecting a strong preference for conferences as the primary channel for academic dissemination. Among the 44 grey literature publications, 41 (93.18%) were blogs, while 3 (6.82%) were white papers, underscoring blogs as the dominant platform for informal knowledge sharing. These trends indicate that chaos engineering, which began as an emerging research area, has evolved into a mature field of interest (Varela-Vaca and Quintero, 2021), attracting significant attention from academia and industry.
@@ 6.2. Number of Sources by Year and Venue Type
Distribution of academic(N = 44) and grey(N = 44) literature studies over years and venue type.
6.3. Publication trends and key Contributors in Academic and Grey Literature
@@ 6.3.1. Academic Literature
First, we analyzed venues based on their name, ranking, and type (journal or conference). As shown in Table 13, prominent venues include key publishers such as IEEE and ACM, which are widely recognized for hosting high-impact conferences and journals in chaos engineering. The venues were categorized using the Computing Research and Education Association of Australasia (CORE) ranking for conferences (CORE 2023) [^3] https://portal.core.edu.au/conf-ranks/ and the Scimago ranking for journals (Scimago 2023) [^4] https://www.scimagojr.com/journalrank.php. Figure 12 presents the distribution of sources in these venues. The data indicates that the largest share of studies is found in unranked venues (12 sources), with a significant number published in Q1 venues reflecting solid academic interest. High-ranking venues, such as A and A*, include several prominent IEEE and ACM conferences, emphasizing the quality of research in the field. Lower-ranked venues, including B, C, and workshops, also contribute to the research landscape, although fewer studies are published in Q2 venues. Table 13 highlights the top 10 primary studies in chaos engineering, ranked by citation count. It should be noted that only the first author of each study is listed, even if there are multiple authors  (Varela-Vaca and Quintero, 2021). Study A42 from Sun Yat-Sen University, China, leads the list, presenting MicroRank, a tool for end-to-end latency issue localization in microservice environments. Following closely are studies A02 and A09 from the Hasso Plattner Institute in Germany, which explore chaos engineering applications in cloud security and continuous auditing and threat detection across multi-cloud environments. KTH Royal Institute of Technology, Sweden, is represented by three influential works, A03, A05, and A06, investigating fault injection and resilience in containerized systems and Java Virtual Machine (JVM) environments. Netflix contributes with study A36, focusing on automating chaos experiments within production environments, while Carnegie Mellon University’s A22 emphasizes service-level fault injection testing via the Filibuster tool. Additional contributions include A26 from Lund University, which presents an implementation framework for chaos engineering, and A35 from the University of Stuttgart, which proposes an approach for prioritizing chaos experiments using established risk analysis techniques. This collection of studies from institutions in countries such as China, Germany, Sweden, and the United States reflects a global commitment to advancing chaos engineering with applications to enhance system resiliency, strengthen cloud infrastructure, and improve reliability in production environments worldwide.
@@ 6.3.2. Grey Literature
Unlike traditional academic literature, which follows a structured format and is ranked using systems such as h-index or Scimago, grey literature lacks standardized organization. It cannot be assessed using typical academic metrics. Instead, alt-metrics (alternative metrics used to measure online attention and engagement), such as comments, likes, shares, and social engagement, are used to gauge the quality and impact of grey literature. Table 14 highlights the top five articles in grey literature based on these alt-metrics. Grey literature on chaos engineering is often published by well-known IT companies (e.g., Microsoft, Gremlin, and Capital One) on company blogs or online publishing platforms (e.g., LinkedIn and Medium). These platforms are essential venues for sharing industry-driven insights in formats such as technical articles, blog posts, and white papers. The alt-metrics used, such as likes, comments, and shares, help measure the influence and engagement of the content within the industry, indicating its impact on chaos engineering. In cases where social engagement data was limited, additional quality assessments were considered, including the reputation of the venues, the authors’ expertise, and the clarity of the content (see Table  3).
As shown in Table 14, the top five grey articles are authored by industry professionals who provide valuable insights into chaos engineering. Key contributors come from prominent organizations, including Lyft, National Australia Bank (NAB), Expedia, Chaos Toolkit, and Eat Takeaway.com. For example, the study G12 from Lyft, published on Medium, received 348 likes and 2 comments for its introduction of a chaos experimentation framework. The study G11 from NAB, also published on Medium, received 156 likes for its focus on enhancing observability in microservices during chaos experiments. Additional notable contributions include the study G15 from Expedia, whose article on Medium earned 128 likes for detailing Kubernetes-integrated chaos engineering frameworks. The study G25, authored on LinkedIn by Chaos Toolkit, received 52 likes and 52 comments, emphasizing the importance of human interaction in chaos experiments. Lastly, the study G14 from Just Eat Takeaway.com, published on Medium, received 24 likes and 1 comment, highlighting the discovery of hidden system vulnerabilities through chaos engineering. These industry contributions have significantly influenced the development of chaos engineering practices, providing practical insights directly from leaders in the field. As discussed earlier, the same quality assessment criteria were applied, focusing on venue reputation, author expertise, and content clarity, when data on social engagement were insufficient (see Table  3).
Table 13. Top-10 of Papers Ordered by Citations in Academic Literature (SC - SCOPUS, GS - Google Scholar).
Rank ID Title Author Affiliation SC GS
1 A42 MicroRank: End-to-End Latency Issue Localization with Extended Spectrum Analysis in Microservice Environments Guangba Yu Sun Yat-Sen University, China 50 72
2 A02 CloudStrike: Chaos Engineering for Security and Resiliency in Cloud Infrastructure Kennedy A. Torkura Hasso Plattner Institute, Germany 39 70
3 A09 Continuous Auditing and Threat Detection in Multi-cloud Infrastructure Kennedy A. Torkura Hasso Plattner Institute, Germany 28 53
4 A03 A Chaos Engineering System for Live Analysis and Falsification of Exception-Handling in the JVM Long Zhang KTH Royal Institute of Technology, Sweden 15 44
5 A05 Observability and Chaos Engineering on System Calls for Containerized Applications in Docker Long Zhang KTH Royal Institute of Technology, Sweden 17 40
6 A36 Automating Chaos Experiments in Production Ali Basiri Netflix in Los Gatos, California. 15 40
7 A26 Getting Started with Chaos Engineering - Design of an Implementation Framework in Practice Hugo Jernberg Lund University and ICA Gruppen AB, Lund/Solna, Sweden 7 21
8 A22 Service-level Fault Injection Testing Christopher S. Meiklejohn Carnegie Mellon University, Pittsburgh, PA, United States 14 19
9 A06 Maximizing Error Injection Realism for Chaos Engineering With System Calls Long Zhang KTH Royal Institute of Technology, Sweden 6 18
10 A35 Identifying and Prioritizing Chaos Experiments by Using Established Risk Analysis Techniques Dominik Kesim University of Stuttgart, Germany 8 16
Table 14. Top-5 Articles Ordered by Altmetrics (Comments / Likes/Claps) in Grey Literature.
Rank ID Comments Likes/Claps Author Affiliation Venue
1 G12 2 348 Kathan Shah & Subhra Mondal Lyft Medium
2 G11 - 156 Prateek Sachan National Australia Bank (NAB) Medium
3 G15 - 128 Nikos Katirzis Expedia Group Medium
4 G25 - 52 Russ Miles Chaos Toolkit Medium
5 G14 1 24 Mathew Moon Just Eat Takeaway.com Medium

### v2

@@ 6.1.4. Deployment
6.2. Chaos Engineering Tools
To further address RQ3.1, we applied the taxonomy developed in this study to evaluate ten widely used chaos engineering tools (T1–T10 in Table 13).
We identified 71 tools mentioned across academic and grey sources during our thematic analysis. We filtered this list by inspecting their GitHub repositories. We excluded tools that had been archived (read-only), were no longer maintained, or lacked sufficient documentation. This screening process reduced the set to 41 active tools with publicly available repositories. From this refined pool of 41 tools, we selected the ten most starred repositories on GitHub as of November 2024. The ‘most stars’ metric was used as a proxy for community interest and real-world relevance  (Borges et al., 2016). Each selected tool was evaluated using our proposed taxonomy, which classifies tools based on four key dimensions: execution environments, automation modes, failure injection strategies, and deployment stages. This mapping is presented in Table 13 and is intended to help engineering teams align tool selection with the architecture of the systems under test(e.g., microservices, containerized environments) and operational goals (e.g.,reducing latency, enhancing scalability) (Kavaler et al., 2019).
For example, T1 and T2 support Kubernetes environments. T1 enables manual testing in both pre-production and production stages, while T2 extends automation with custom code and integration, targeting pre-production use. T3 to T5 offer a broader platform support, including Kubernetes, Docker, and VMs, and provide full automation features such as workflow integration, prioritization, and custom code, applicable across all deployment stages. T6 and T7 focus on Kubernetes and containers. Both support manual, semi-automated, and fully automated modes; T6 emphasizes hybrid setups with prioritization, while T7 supports Docker and targets production-grade testing. T8 and T9 operate across Kubernetes and virtual machines. T8 supports complete automation and is usable in both staging and live systems, while T9 focuses on pre-production with advanced integration options. Finally, T10 is tailored for hybrid cloud environments, supporting Kubernetes and providing full automation across both pre-production and production environments.
Table 13. Top-10 Most Used Chaos Engineering Platforms.
Docker Kubernetes
Virtual
Machine
Serverless
Function
Private Public Hybrid
T1 Chaos Monkey
T2 Toxiproxy
T3 Chaos Mesh
T4 ChaosBlade
T5 LitmusChaos
T6 Kube-monkey
T7 Pumba
T8 Chaos Toolkit
T9 Powerfulseal
T10 Chaoskube
6.3. Adoption Practices (RQ3.2)
To address RQ3.2, we highlight best practices and common pitfalls in adopting chaos engineering derived from our literature sources. Successfully adopting chaos engineering requires a structured approach that balances best practices with avoiding common mistakes. To minimize risk during experimentation, organizations should follow well-defined guidelines for planning, executing, and evaluating chaos experiments  (Siwach et al., 2022; Green, 2023). Table 14 summarizes essential best practices, such as starting with small, controlled experiments, setting clear objectives, automating tests, and closely monitoring system behavior. Additionally, organizations should limit the scope of experiments, encourage cross-team collaboration, and continuously refine processes based on experiment outcomes. While following best practices is crucial, organizations must also avoid common pitfalls, such as those outlined in Table 15, which include failing to tailor experiments to specific systems, neglecting proper planning and hypothesis formation, and lacking adequate safety mechanisms. By adhering to best practices and avoiding these common pitfalls, organizations can confidently integrate chaos engineering into their systems and achieve successful outcomes (Siwach et al., 2022). For full details, see the code-to-theme mapping in the online appendix (Section 3.6).
Table 14. Best Practices.
No Best Practices Description Articles Total
1 Start Small and Scale Up Start with manageable tests, then expand gradually. (Fogli et al., 2023; Torkura et al., 2020; Bremmers, 2021; Patel, 2022; Gill, 2021; Chen et al., 2022; Siwach et al., 2022; Jernberg et al., 2020; Anwar and Arunachalam, 2019; Palani and Gupta, 2023a; Kadikar, 2023; Green, 2023) 12
6 Document and Improve Record experiments to continually improve the system. (Torkura et al., 2020; Zhang et al., 2019; Inc., 2023; Bremmers, 2021; Patel, 2022; Gill, 2021; Siwach et al., 2022; Palani and Gupta, 2023a; Kadikar, 2023; Green, 2023; Kesim et al., 2020) 11
Table 15. Bad Practices.
No Bad Practices Description Articles Total
1 Misaligned Chaos Chaos experiments that don’t match the system yield misleading results. (Fogli et al., 2023; Torkura et al., 2020; Zhang et al., 2019; Dedousis et al., 2023; Vu et al., 2022; Nikolaidis et al., 2023; Ma et al., 2023; Sondhi et al., 2021; Jernberg et al., 2020; Konstantinou et al., 2021; Park et al., 2019; Zhang et al., 2023; Rivera et al., 2023; Patel, 2022; Le, 2022; Lardo, 2019; Hochstein and Rosenthal, 2016; Basiri et al., 2016) 18
2 Inadequate Experimentation Without planning, clear goals, and baseline tracking, assessing chaos test impacts becomes difficult. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Jakkaraju, 2020; Le, 2022; Hawkins, 2020; Anwar and Arunachalam, 2019; Dua, 2024; National Australia Bank, 2020; Mooney, 2023; Starr et al., 2022; Palani and Gupta, 2023a; Gill, 2022; Fawcett, 2020; Gianchandani et al., 2022; Bairyev, 2023; Abdul, 2024; Azure Readiness, 2023; Kadikar, 2023; Green, 2023; Aharon, 2024; Krivas and Portela, 2020; Kostic, 2024; Frank et al., 2021b; Jernberg et al., 2020; Konstantinou et al., 2021; Ji et al., 2023; Soldani and Brogi, 2021; Park et al., 2019; Wachtel, 2023; Alvaro and Tymon, 2017; Sousa et al., 2018) 33
3 Safety Mechanism Oversight Not testing rollback and safety features risks uncontrolled failures. (Inc., 2023; Wickramasinghe, 2023; Gill, 2021; Bremmers, 2021; Patel, 2022; Jakkaraju, 2020; Bhaskar, 2022; Katirtzis, 2022; Dua, 2024; Starr et al., 2022; Tivey et al., 2019; Palani and Gupta, 2023a; Bairyev, 2023; Sbiai et al., 2023; Kadikar, 2023; Green, 2023; Vinisky, 2024; Pierce et al., 2021; Al-Said Ahmad et al., 2024; Vu et al., 2022; Ikeuchi et al., 2020; Chen et al., 2022; Naqvi et al., 2022; Soldani and Brogi, 2021; Konstantinou et al., 2021; Alvaro and Tymon, 2017; Sousa et al., 2018) 27
6.4. Evaluation Approaches (RQ3.3)
@@ 6.2. Chaos Engineering Tools
Evaluation refers to determining the effectiveness of chaos engineering experiments by measuring their impact on system stability, the system’s ability to handle induced failures, and the growth and maturity of the organization’s chaos engineering practices. To address RQ3.3, we categorized evaluation approaches into quantitative and qualitative methods based on themes derived from our systematic coding and thematic analysis (see Section 3.5 and detailed mappings in the online appendix 3.6).
@@
  • Business Impact Evaluation assesses how chaos experiments affect key business metrics, including revenue, customer churn, and operational efficiency (Inc., 2023; Jernberg et al., 2020). This involves understanding specific financial losses, customer complaints, and service-level breaches resulting from experiments(National Australia Bank, 2020; Palani and Gupta, 2023a). For instance, a chaos experiment that causes downtime during a product launch could result in significant lost sales and a spike in customer service tickets (Ji et al., 2023; Starr et al., 2022). This evaluation is critical to linking technical failures with direct business consequences, enabling more targeted investment in stability measures where it matters most (Manna, 2021; Moon, 2022).
@@ 7. Current State of Chaos Engineering (RQ4)
This section addresses RQ4 by exploring how chaos engineering has evolved in both academic and industry contexts. Specifically, it first presents the classification of research in chaos engineering, then examines the number of sources published per year and venue type (RQ4.1), and finally reviews the venue rankings and key contributors in the field (RQ4.2).
7.1. Number of Sources by Research Type (RQ4.1)
@@
To better understand the current state of chaos engineering research (addressing RQ4.1), we classified the primary studies based on the research classification framework proposed by Wieringa et al. (Wieringa et al., 2006). This framework includes the following categories:
  • Evaluation research: Investigates real-world problems through empirical methods such as case studies and surveys.
  • Philosophical papers: Propose new conceptual models, frameworks, or perspectives.
  • Solution proposals: Present novel techniques or tools, typically without full empirical validation.
  • Validation research: Rigorously assess proposed solutions in controlled environments (e.g., simulations, experiments).
  • Opinion papers: Express the author’s viewpoint on a topic, often arguing for or against certain practices.
  • Experience reports: Describe applied practices and lessons from real-world projects.
@@
The first author did the classification, which was then reviewed by the second author, with all discrepancies resolved through discussion. The results of this classification, shown in Figure 11, indicate that solution proposal papers were the most frequent type of research between 2016 and 2024. In the academic literature, solution proposal papers were the most prevalent, with nine papers in 2021 and 2022, and slightly fewer (eight) in 2023. On the other hand, the grey literature showed a slight upward trend in experience reports, rising from six in 2022 to eight in 2023. This classification highlights the evolving distribution and focus of research activities, emphasizing the dominance of solution proposals in academic literature and the increasing role of practical experience reports in the grey literature.
@@
Distribution of academic(N = 49) and grey(N = 47) literature studies over years and venue type.
7.2. Number of Sources by Year and Venue Type
To further address RQ4.1, we analyzed the yearly distribution of academic and grey literature on chaos engineering from 2016 to early 2024 (see Figure 9). Academic output was limited from 2016 to 2018 but grew significantly from 2019, peaking in 2021 and 2022 with 13 and 12 studies, respectively, mainly in the form of conference papers. Grey literature remained minimal until 2019, rising steadily to a peak of 12 sources in 2023, dominated by blog posts and a few white papers. A slight drop in 2024 is likely due to the incomplete year (Werbińska-Wojciechowska and Winiarska, 2023). Overall, these trends underscore the growing scholarly and industrial interest in chaos engineering. Further breakdowns of publication types, venues, and source classifications are available in our online appendix (Section 3.6).
7.3. Publication trends and key Contributors in Academic and Grey Literature (RQ4.2)
@@ 7.3.1. Academic Literature
First, we analyzed publication venues based on their names, types (journal or conference), and rankings. The venues were categorized using the Computing Research and Education Association of Australasia (CORE) ranking for conferences (CORE 2023) [^4] https://portal.core.edu.au/conf-ranks/ and the Scimago ranking for journals (Scimago 2023) [^5] https://www.scimagojr.com/journalrank.php. As shown in Figure 11, the largest share of studies appeared in unranked venues (12), though many were published in top-tier sources—particularly CORE A/A* conferences and Scimago Q1 journals, often from IEEE and ACM. This suggests increasing scholarly validation of chaos engineering research. Lower-ranked venues (B, C) and workshops contributed modestly, with few studies published in Q2 venues. To highlight influential contributions, we identified the top 10 academic studies based on citation counts from Scopus and Google Scholar (first author only). The leading study, authored by Netflix (Basiri et al., 2016), has become foundational to the field, followed by diverse works from institutions in China, Germany, Sweden, and the U.S. that span topics such as end-to-end latency localization (e.g., MicroRank) (Yu et al., 2021), chaos-based security testing (Torkura et al., 2020), and exception handling in JVM environments (Zhang et al., 2019). These publications reflect a global effort toward improving system resiliency and reliability through chaos engineering. Full details are provided in our online appendix (Section 3.6).
@@ 7.3.2. Grey Literature
Unlike academic publications, grey literature lacks formal rankings and peer review. To evaluate its quality and impact, we utilized alternative metrics (e.g., likes, comments, shares) and, when social engagement data were limited, applied qualitative criteria such as venue reputation, author expertise, and content clarity (see Table 5). The most engaged grey literature source was “Chaos Monkey Upgraded” by Netflix  (Hochstein and Rosenthal, 2016), which received 443 likes and 2 comments. Other widely read posts include Lyft’s framework introduction (348 likes) (Shah, 2021), NAB’s work on observability (156 likes) (Mondal and Sachan, 2020), and technical contributions from Expedia (Katirtzis, 2022) and the Chaos Toolkit community (Miles, 2019a). These practitioner-authored works offer valuable, experience-based perspectives that complement academic research. Further details, including top-ranked articles and alt-metric data, are available in the online appendix (Section 3.6).
[…]
Research Type